Marc Horowitz 24.151 March 6, 1992

I

(1) True.

(2) True.

(3) False. Senses are not concrete objects, so the sese of ‘the morning star’ cannot be a planet, which is a concrete object.

(4) False. The image associated with ‘the morning star’ is that of a point of light in the sky.

(5) True.

(6) False. Senses express references. ‘ ‘The morning star’ ’ is a sign, not a sense.

(7) False. Again, a sense expresses a reference. Senses do not present references.

II

No. For example, ‘I said “My friend visited me” ’, and ‘I said “I was visited by my frienf” ’ have different truth values, although the statements have the same sense.

III

No, they are not the same. ‘the morning star’ and ‘the sense of ‘the morning star’,’ are not the same. Senses are not concrete objects. (I get the feeling I’m really missing something here, or you did not ask the question you meant to ask.)

This sentence does not present a problem for Frege. It is a simple truth-value sentence, with two parts. For the first part to be true, the reference of ‘the morning star’ must be a planet. For the second part to be true, John must be aware of this fact.